A multimode step index fiber with a core diameter of 80 μm and a relative index difference of 1.5 % is

operating at a wavelength of 0.85 μm. If the core refractive index is 1.48, estimate the normalized

frequency for the fiber and number of guided modes.


1
Expert's answer
2021-03-09T15:34:42-0500

Answer

Here wavelegth

"\lambda=085\mu m"

Radius of core

"a=D/2=40\mu m"

"n_1=" core refractive index=1.48

"\Delta=0.015"

Putting all value in below equation

normalised frequency of fiber

"f= (2π/λ) a n_1 (2Δ)^{0.5}\\=\frac{2\times3.14\times40\mu m\times1.48}{0.085\mu m}\times(2\times0.015)^{0.5}\\= 75.8Hz"

Now new parameters

"V=\frac{2\pi a}{\lambda}(\sqrt{n_1^2-n_2^2})\\= \frac{2\times3.14\times0.00004}{0.85\mu m}(\sqrt{0.015}) \\ =75.80"

the number of guided modes.

"M ≈\frac{ V^2} { 2 }= 2873"

(i.e. nearly 3000 guided modes!)